OpenStudy (anonymous):

please help me, Convert to rectangular coordinates. Use exact values. (4sqr3,-pi/4)

OpenStudy (mathmale):

The given point is in which Quadrant? How do you know? What is the length of the ray from the origin to the point?

OpenStudy (anonymous):

it doesn't say anything else...

OpenStudy (theeric):

\(x=r\cos\theta\) \(y=r\sin\theta\) It is that easy :) 

OpenStudy (mathmale):

You are given (4sqr3,-pi/4). These are the polar coordinates of a point in the xy plane. Such a point is specified by (r, theta), where r is the distance of the point from the origin and theta is the angle that that distance (ray) makes with the positive x-axis. So, in this case, the distance of your point from the origin is 4Sqrt(3) (about 4*1.732), and this distance is at an angle of pi/4 clockwise, as measured from the positive x-axis.